# SecureFalcon: Are We There Yet in Automated Software Vulnerability Detection with LLMs?

### v2

@@ I Introduction
As we are undoubtedly passing through a digital age, technology affects every aspect of our lives [1]. In such a context, vulnerability detection tools are essential safeguards in our continuously evolving digital landscape [2, 3]. With the emergence of new technologies, the palette of cyber threats becomes wider and more sophisticated in terms of employed techniques. Vulnerability detection tools assist in scanning, probing, and inspecting software, identifying weaknesses that could serve as entry points for potential malicious users or attackers. While integral to software security measures, vulnerability detection tools encounter inherent limitations in their scope and efficiency. Such tools typically depend on known patterns and signatures derived from synthetic datasets, making them less effective for detecting bugs in real-world software [4]. Furthermore, the datasets used by these tools are often too small [5] or have a skewed distribution of vulnerable programs [6], hampering the accuracy of static analyzer tools. Although the advancements in deep learning (DL) seem promising for vulnerability detection [7, 8, 9, 10], their accuracy heavily relies on the data quality as well [6, 11]. Most of the popular datasets are either fully synthetic [12, 13], non-compilable [14, 15, 16], or have an unfair distribution of vulnerable vs non-vulnerable code [11, 15, 17]. Moreover, the labels associated with these datasets are subject to the detection method used. For instance, manual labeling is affected by human errors, whereas using static analysis tools to label the data can result in high false positives [18, 19]. While the DL-based approaches are often employed for quick inference, they may compromise the model’s accuracy in detecting bugs. Formal Verification (FV) approaches such as Model Checking (MC) are preferred for verification in safety-critical systems [20]. Even though MC provides safety assurances for software systems [21, 22], they are often expensive, even for a small piece of code. Bounded Model Checking (BMC) [23, 24] improves the traditional MC techniques by restricting the exploration to a predefined bound or depth. To prove safety in BMC for programs, we must compute the completeness threshold (CT), which can be smaller than or equal to the maximum number of loop iterations occurring in the program. Although BMC offers performance improvement to some extent, it is still expensive.
@@
The red-highlighted area represents the gap that can only be closed through rigorous mathematical formal verification [26]. However, overall, formal program verification is undecidable [27, 28]. We cannot devise a computational method to decide whether an arbitrary program is error-free. Much progress has been made in program verification. Nowadays, verification methods create abstractions to determine whether the program is error-free. Still, for an arbitrary program that includes unbounded memory usage, we cannot give $100$% certainty that it is error-free due to the well-known halting problem [27]. However, to achieve high accuracy detection rate for vulnerable code almost instantly with a fine-tuned model would be a significant achievement. We aim to minimize this gap (red-highlighted area) as much as possible and dramatically reduce vulnerability detection times, making it practical for real-time code completion frameworks. While it is unrealistic to expect LLMs to match the results of BMC verification tools at this stage, there is promise in using ML techniques to swiftly identify the most vulnerable programs. Integrating a robust BMC tool with a pre-trained LLM can enhance the pace of vulnerability detection, benefiting the overall software development process. The success of a vulnerability detection model greatly depends on the quality and relevance of the dataset it uses. Thus, a dataset that provides a balanced distribution of vulnerable and non-vulnerable programs, mirrors real-world coding, and is labeled with a reliable method can enhance the model’s accuracy. In contrast to most available C/C++ vulnerability datasets, the FormAI dataset [29] 1 has high labeling accuracy as it is labeled by an unbounded model checker ESBMC [24], 2 contains compilable programs, 3 evenly distributes vulnerable and neutral programs, and 4 is generated by LLM trained on real-world open source software, thus, mimicking the prominent developer errors and making it a highly suitable dataset for training a vulnerability detection model. The original contributions of our work are summarized as follows:
@@
  • We introduce SecureFalcon, a lightweight and innovative LLM model with only $121$ million parameters, built on the foundational 40B-parameter FalconLLM architecture. This model offers significant enhancements and optimizations specifically tailored for security analysis. SecureFalcon underwent fine-tuning with the introduction of the FormAI dataset, a specialized collection designed to accurately classify vulnerabilities within C/C++ code samples. This process was supported by using the Efficient SMT-based Context-Bounded Model Checker (ESBMC) tool, enhancing the model’s detection capabilities.
@@
  • Since the formal verification tools cannot detect all Common Weakness Enumerations (CWEs), such as CWE-78, we created an aggregated dataset called FalconVulnDB, incorporating the SySeVR framework, Draper VDISC, Bigvul, Diversevul, SARD Juliet, and ReVeal datasets, to enrich SecureFalcon’s training by leveraging an extensive compilation of public datasets. These resources collectively cover examples of the top $25$ most critical software weaknesses identified by the Common Weakness Enumeration (CWE), including but not limited to CWE-119, CWE-120, CWE-476, CWE-122, CWE-190, CWE-121, CWE-78, CWE-787, CWE-20, and CWE-762. The FalconVulnDB dataset was used to compensate for the lack of real project data in the FormAI dataset.
@@ II Motivation
where $I$ is the initial states set, and $T(s_{i},s_{i+1})$ the transition relation of $\mathcal{TS}$. This formula captures $\mathcal{TS}$ executions of length $k$, and if it is satisfied, a state violates $\phi$ within the $k$ limit. A counterexample, or trace, for a violated $\phi$ is a sequence of states $s_{1},\ldots,s_{k}$. If Equation (1) is unsatisfiable, no error state exists within $k$ steps, indicating no vulnerabilities in the program up to the $k$ bound. Checking property violations using BMC techniques can be time-consuming even for relatively small programs due to lengthy loops that require unwinding or intricate function calls [24]. Consider Listing 1 where an arithmetic overflow occurs due to the nature of the nested loops and the accumulation of values in the sum variable. The code consists of two nested loops iterating from 00 to $9999,9$ 999 (i and j both range from 00 to $999,9$ 999). Within each iteration, the product of i * j is added to the sum variable. Since the sum variable accumulates these products (sum += i * j;), it is possible for the sum to exceed the maximum value represented by an int. When this happens, it causes arithmetic overflow, which means the result is beyond the range represented by the data type, resulting in an incorrect value being stored in the sum variable.
@@ IV-B SecureFalcon Model Architecture
The SecureFalcon model, derived from the 40B parameter FalconLLM, includes components shown in Fig. 2, featuring out_features=12 for multi-classification and out_features=2 for binary classification.
@@ V-1 FormAI dataset
SecureFalcon uses the FormAI dataset [^4] https://ieee-dataport.org/documents/formai-dataset-large-collection-ai-generated-c-programs-and-their-vulnerability[29] for fine-tuning the model. The FormAI dataset includes $112,000$ compilable C code snippets created using the GPT-3.5-turbo model through a dynamic zero-shot prompting method. Since this GPT model from OpenAI [^5] https://openai.com/ is trained using real-world open-source software repositories, the code it generates closely mimics real-world code behavior. The produced C samples are subsequently verified using the Efficient SMT-based Context-Bounded Model Checker (ESBMC) [gadelha2018esbmc]. ESBMC is set with a verification timeout of 30 seconds per sample. According to the 2023 SV-COMP results [^6] https://sv-comp.sosy-lab.org/2023/results/results-verified/quantilePlot-Overall.svg, this verifier has successfully addressed the most significant number of verification tasks. The generated C samples are classified into three categories: verification successful, verification unknown, and verification failed. The verification failed class is further divided into specific vulnerability types: arithmetic overflow, buffer overflow, array bounds violation, NULL pointer dereference, division by zero, and others. The dataset provides insights into the distribution of vulnerabilities and serves as a valuable resource for vulnerability analysis and testing.
@@
V-2 FalconVulnDB Dataset
@@ VI-B Experimental Results
Figure 6: Confusion matrix of SecureFalcon 121M/44M classification using FormAI dataset.
@@
VI-C Comparison of SecureFalcon

### v3

@@ I Introduction
As we are undoubtedly passing through a digital age, technology affects every aspect of our lives [1]. In such a context, vulnerability detection tools are essential safeguards in our continuously evolving digital landscape [2]. With the emergence of new technologies, the palette of cyber threats becomes wider and more sophisticated in terms of employed techniques. Vulnerability detection tools assist in scanning, probing, and inspecting software, identifying weaknesses that could serve as entry points for potential malicious users or attackers. While integral to software security measures, vulnerability detection tools encounter inherent limitations in their scope and efficiency. Such tools typically depend on known patterns and signatures derived from synthetic datasets, making them less effective for detecting bugs in real-world software [3]. Furthermore, the datasets used by these tools are often too small [4] or have a skewed distribution of vulnerable programs [5], hampering the accuracy of static analyzer tools. Although the advancements in deep learning (DL) seem promising for vulnerability detection [6, 7, 8, 9], their accuracy heavily relies on the data quality as well [5, 10]. Most of the popular datasets are either fully synthetic [11, 12], non-compilable [13, 14, 15], or have an unfair distribution of vulnerable vs non-vulnerable code [10, 14, 16]. Moreover, the labels associated with these datasets are subject to the detection method used. For instance, manual labeling is affected by human errors, whereas using static analysis tools to label the data can result in high false positives [17, 18]. While the DL-based approaches are often employed for quick inference, they may compromise the model’s accuracy in detecting bugs. Formal Verification (FV) approaches such as Model Checking (MC) are preferred for verification in safety-critical systems [19]. Even though MC provides safety assurances for software systems [20, 21], they are often expensive, even for a small piece of code. Bounded Model Checking (BMC) [22, 23] improves the traditional MC techniques by restricting the exploration to a predefined bound or depth. To prove safety in BMC for programs, we must compute the completeness threshold (CT), which can be smaller than or equal to the maximum number of loop iterations occurring in the program. Although BMC offers performance improvement to some extent, it is still expensive.
@@
The red-highlighted area represents the gap that can only be closed through rigorous mathematical formal verification [25]. However, overall, formal program verification is undecidable [26, 27]. We cannot devise a computational method to decide whether an arbitrary program is error-free. Much progress has been made in program verification. Nowadays, verification methods create abstractions to determine whether the program is error-free. Still, for an arbitrary program that includes unbounded memory usage, we cannot give $100$% certainty that it is error-free due to the well-known halting problem [26]. However, to achieve high accuracy detection rate for vulnerable code almost instantly with a fine-tuned model would be a significant achievement. We aim to minimize this gap (red-highlighted area) as much as possible and dramatically reduce vulnerability detection times, making it practical for real-time code completion frameworks. While it is unrealistic to expect LLMs to match the results of BMC verification tools at this stage, there is promise in using ML techniques to swiftly identify the most vulnerable programs. Integrating a robust BMC tool with a pre-trained LLM can enhance the pace of vulnerability detection, benefiting the overall software development process. The success of a vulnerability detection model greatly depends on the quality and relevance of the dataset it uses. Thus, a dataset that provides a balanced distribution of vulnerable and non-vulnerable programs, mirrors real-world coding, and is labeled with a reliable method can enhance the model’s accuracy. In contrast to most available C/C++ vulnerability datasets, the FormAI dataset [28] 1 has high labeling accuracy as it is labeled by the Efficient SMT-based Context-Bounded Model Checker (ESBMC) [23], 2 contains compilable programs, 3 evenly distributes vulnerable and neutral programs, and 4 is generated by LLM trained on real-world open source software, thus, mimicking the prominent developer errors and making it a highly suitable dataset for training a vulnerability detection model. The original contributions of our work are summarized as follows:
@@
  • We introduce SecureFalcon, a lightweight and innovative LLM model with only $121$ million parameters, built on the foundational 40B-parameter FalconLLM architecture. This model offers significant enhancements and optimizations specifically tailored for security analysis. SecureFalcon was enhanced through fine-tuning using the FormAI dataset, a specialized collection designed to accurately classify vulnerabilities in C code samples. This process was supported by using the ESBMC tool, enhancing the model’s detection capabilities.
  • To address the limitations of formal verification tools, which are highly effective at detecting memory-related issues such as buffer overflows (CWE-121, CWE-122) and array bound violations (CWE-129) but are unable to identify vulnerabilities like OS command injection (CWE-78) or SQL injection (CWE-89), we created an aggregated dataset named FalconVulnDB. This dataset integrates resources from the SySeVR framework, Draper VDISC, Bigvul, Diversevul, SARD Juliet, and ReVeal datasets to enhance SecureFalcon’s training with a comprehensive collection of public datasets.
@@
    These resources collectively cover examples of the top $25$ most critical software weaknesses identified by the Common Weakness Enumeration (CWE), including but not limited to CWE-119, CWE-120, CWE-476, CWE-122, CWE-190, CWE-121, CWE-78, CWE-787, CWE-20, and CWE-762. The FalconVulnDB dataset was used to compensate for the lack of real project data in the FormAI dataset.
@@ II Motivation
where $I$ is the initial states set, and $T(s_{i},s_{i+1})$ the transition relation of $\mathcal{TS}$. This formula captures $\mathcal{TS}$ executions of length $k$, and if it is satisfied, a state violates $\phi$ within the $k$ limit. A counterexample, or trace, for a violated $\phi$ is a sequence of states $s_{1},\ldots,s_{k}$. If Equation (1) is unsatisfiable, no error state exists within $k$ steps, indicating no vulnerabilities in the program up to the $k$ bound. Checking property violations using BMC techniques can be time-consuming even for relatively small programs due to lengthy loops that require unwinding or intricate function calls [23]. Consider Listing 1 where an arithmetic overflow occurs due to the nature of the nested loops and the accumulation of values in the sum variable. The code consists of two nested loops iterating from 00 to $9999,9$ 999. Within each iteration, the product of i * j is added to the sum variable. Since the sum variable accumulates these products (sum += i * j;), the sum eventually surpasses the maximum value that an int can represent. When this happens, it causes arithmetic overflow, which means the result is beyond the range represented by the data type, resulting in an incorrect value being stored in the sum variable.
@@ IV-B SecureFalcon Model Architecture
The SecureFalcon model, derived from the 40B parameter FalconLLM, includes components shown in Fig. 2, with out_features=12 for multi-class classification and out_features=2 for binary classification.
@@ V-1 FormAI dataset
SecureFalcon uses the FormAI dataset[28] for fine-tuning the model. The FormAI dataset includes $112,000$ compilable C code snippets created using the GPT-3.5-turbo model through a dynamic zero-shot prompting method. Since this GPT model from OpenAI [^4] https://openai.com/ is trained using real-world open-source software repositories, the code it generates closely mimics real-world code behavior. The produced C samples are subsequently verified using ESBMC [62].
ESBMC is set with a verification timeout of 30 seconds per sample. According to the 2023 SV-COMP results [^5] https://sv-comp.sosy-lab.org/2023/results/results-verified/quantilePlot-Overall.svg, this verifier has successfully addressed the most significant number of verification tasks. The generated C samples are classified into three categories: verification successful, verification unknown, and verification failed. The verification failed class is further divided into specific vulnerability types: arithmetic overflow, buffer overflow, array bounds violation, NULL pointer dereference, division by zero, and others. The dataset provides insights into the distribution of vulnerabilities and serves as a valuable resource for vulnerability analysis and testing.
@@
V-2 Limitations of AI-Generated Training Data
While including the FormAI dataset allowed us to scale our training corpus with synthetically generated samples rapidly, we recognize that this dataset alone may not fully capture the complexities of real-world code. AI-generated code often lacks the historical context, unconventional patterns, and domain-specific styles in human-written codebases. To mitigate this limitation, we integrated the SySeVR framework, Draper VDISC, Bigvul, Diversevul, SARD Juliet, and ReVeal datasets into a consolidated resource called FalconVulnDB. This aggregated dataset encompasses examples of the top 25 most critical software weaknesses (CWE), such as CWE-119, CWE-120, CWE-476, CWE-122, CWE-190, CWE-121, CWE-78, CWE-787, CWE-20, and CWE-762, drawn from public, human-curated sources.
By leveraging FalconVulnDB, we aim to compensate for the scarcity of genuine project data in the FormAI corpus. Unlike AI-generated code, these publicly available datasets capture various coding habits, legacy structures, and context-driven vulnerabilities. The combination of AI-synthesized and human-sourced data fosters a more balanced training environment, encouraging our model to learn from the syntactic consistency of machine-generated samples and the gritty realism of authentic software repositories.
V-3 FalconVulnDB Dataset
@@
V-4 Data Integration, Standardization, and Bias Mitigation
We acknowledge the known data quality issues associated with widely used datasets like BigVul, as highlighted in prior studies such as LineVul [63]. To ensure the reliability of our results, we implemented a rigorous pre-processing pipeline that included deduplication to eliminate redundant records, validation of vulnerability labels by cross-referencing with publicly available vulnerability descriptions, and removing incomplete or ambiguous entries.
We employed a structured integration and standardization process to ensure consistency, reliability, and transparency in combining these heterogeneous sources. First, we established a unified schema that standardized fields such as vulnerability labels, CWE tags, and severity metrics. For datasets that provided direct CWE mappings, we incorporated them as-is; for those offering only CVE references or binary vulnerability labels (e.g., ReVeal), we cross-checked these identifiers against authoritative databases (e.g., NVD) to map or infer the most likely CWE categories whenever possible. For instance, if a function marked as vulnerable in ReVeal referenced a CVE later confirmed by NVD as related to a buffer overflow vulnerability, we annotated that sample with the corresponding CWE (e.g., CWE-120). When no reliable CWE classification could be assigned, we preserved the binary vulnerability label to maintain dataset diversity.
We also addressed overlaps and conflicts among datasets. In one example, the same function appeared in both Big-Vul and SySeVR, but SySeVR indicated it had been patched, while Big-Vul still listed it as vulnerable. We adopted a hierarchical decision approach in such cases: we consulted NVD and official vendor advisories to verify the patch status. If confirmed, the function was recorded as patched and no longer considered an active vulnerability. Another example involved duplicates in which multiple datasets reported the same CVE but provided slightly different severity scores. Here, we normalized the severity by referencing a standard scale (e.g., CVSS) and took the median of reported severity metrics, ensuring a more robust and standardized representation of the vulnerability’s impact.
To reduce bias, we examined the data distribution across various dimensions—such as vulnerability type, programming language, and software domain—and identified areas of overrepresentation. For example, if stack-based buffer overflows (CWE-121) were disproportionately frequent due to their abundance in a particular dataset, we incorporated more diverse examples from other sources or filtered out duplicates to achieve a more balanced distribution. Additionally, we validated portions of FalconVulnDB against established benchmarks like NVD. For example, if our integrated dataset suggested null pointer dereferences (CWE-476) were exceedingly rare compared to NVD distributions, we revisited our filtering and inclusion criteria to better align with real-world data.
@@ VI-B Experimental Results
Figure 6: Confusion matrix of SecureFalcon 121M classification using FormAI dataset.
@@
As evident in the matrix, the model demonstrates strong performance in detecting the "Not-Vulnerable" class, with a high number of true positives and minimal misclassifications. However, for certain challenging vulnerability types, such as CWE-20 (Improper Input Validation) and CWE-787 (Out-of-Bounds Write), the model struggles to achieve the same level of precision and recall. This is likely due to the limited support for these classes (440 and 524 samples, respectively) in the dataset, leading to a higher rate of misclassification into more frequent categories, such as CWE-119 (Memory Corruption) or the "Other" class.
Furthermore, rare and complex classes like CWE-476 (NULL Pointer Dereference) and CWE-190 (Integer Overflow or Wraparound) show lower performance compared to well-represented classes, such as CWE-78 (OS Command Injection) and CWE-121 (Stack-Based Buffer Overflow), which have higher support and more distinct patterns. The confusion matrix highlights these discrepancies, offering valuable insights into the model’s misclassification patterns. For future work, we plan to enhance the dataset by augmenting samples for underrepresented classes and leveraging advanced techniques such as the focal loss [64] to prioritize learning on these challenging categories. These improvements will help address the current limitations and improve the model’s robustness for detecting rare vulnerabilities.
@@
TABLE XII: Comprehensive Evaluation Metrics for SecureFalcon 121M with LR = 2e-5 using the FalconVulnDB Dataset.
Metric Value
Cohen’s Kappa 0.86
MCC (Matthews Corr. Coeff.) 0.79
Macro F1-Score 0.74
Weighted F1-Score 0.90
ROC-AUC (Macro Avg) 0.88
PR-AUC (Macro Avg) 0.85
Accuracy 0.92
Macro Precision 0.79
Macro Recall 0.72
Weighted Precision 0.91
Weighted Recall 0.91
Specificity (Macro Avg) 0.89
Log Loss 0.35
Brier Score 0.15
Hamming Loss 0.08
VI-C Comprehensive Evaluation Metrics
Table XII presents a comprehensive evaluation of the SecureFalcon 121M model trained with a learning rate of $2\times 10^{-5}$ on the FalconVulnDB dataset, showcasing robust performance across multiple metrics. The following metrics are used in our analysis:
  • Cohen’s Kappa: Measures the agreement between the predicted and actual classifications, accounting for the possibility of agreement occurring by chance.
    $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}$ (11)
    where $P_{o}$ is the observed agreement and $P_{e}$ is the expected agreement by chance.
  • MCC (Matthews Correlation Coefficient): Evaluates the quality of binary classifications, considering true and false positives and negatives.
    $\text{MCC}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (12)
    where $TP$ (True Positive) and $TN$ (True Negative) represent the correctly predicted positive and negative instances, respectively, while $FP$ (False Positive) and $FN$ (False Negative) denote the incorrectly predicted positive and negative instances.
  • Macro F1-Score: The unweighted average of F1-scores for each class, treating all classes equally.
[…]
    where $C$ is the number of classes, $P_{i}$ is the precision, and $R_{i}$ is the recall for class $i$.
  • Weighted F1-Score: The average of F1-scores weighted by the number of true instances for each class.
[…]
    where $w_{i}=\frac{\text{Support}_{i}}{\sum_{j=1}^{C}\text{Support}_{j}}$.
  • ROC-AUC (Macro Avg): The average Area Under the Receiver Operating Characteristic Curve across all classes, assessing the model’s ability to distinguish between classes.
[…]
  • PR-AUC (Macro Avg): The average Area Under the Precision-Recall Curve across all classes, focusing on the trade-off between precision and recall.
[…]
  • Macro Precision: The unweighted average precision across all classes, emphasizing the model’s ability to correctly identify positive instances.
    $\text{Precision}_{\text{macro}}=\frac{1}{C}\sum_{i=1}^{C}P_{i}$ (17)
  • Macro Recall: The unweighted average recall across all classes, highlighting the model’s ability to capture all relevant instances.
    $\text{Recall}_{\text{macro}}=\frac{1}{C}\sum_{i=1}^{C}R_{i}$ (18)
  • Weighted Precision: The precision score weighted by the number of true instances for each class, reflecting overall precision performance.
    $\text{Precision}_{\text{weighted}}=\sum_{i=1}^{C}w_{i}\times P_{i}$ (19)
  • Weighted Recall: The recall score weighted by the number of true instances for each class, indicating overall recall performance.
    $\text{Recall}_{\text{weighted}}=\sum_{i=1}^{C}w_{i}\times R_{i}$ (20)
  • Specificity (Macro Avg): The average specificity across all classes, measuring the model’s ability to correctly identify negative instances.
    $\text{Specificity}_{\text{macro}}=\frac{1}{C}\sum_{i=1}^{C}\frac{TN_{i}}{TN_{i}+FP_{i}}$ (21)
  • Log Loss: Evaluates the uncertainty of the model’s predictions based on the probability estimates.
    Log $\text{Log Loss}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right]$ (22)
    where $N$ is the number of samples, $y_{i}$ is the true label, and $p_{i}$ is the predicted probability.
  • Brier Score: Measures the mean squared difference between predicted probabilities and actual outcomes.
    Brier $\text{Brier Score}=\frac{1}{N}\sum_{i=1}^{N}(p_{i}-y_{i})^{2}$ (23)
  • Hamming Loss: The fraction of incorrect labels to the total number of labels.
    Hamming $\text{Hamming Loss}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(y_{i}\neq\hat{y}_{i})$ (24)
A high Accuracy of 0.92 and a Weighted F1-Score of 0.90 suggest strong overall effectiveness and good handling of class imbalances. Cohen’s Kappa of 0.86 and MCC of 0.79 underscore the reliability and stability of the model’s predictions. At the same time, the ROC-AUC (Macro Avg) of 0.88 and PR-AUC (Macro Avg) of 0.85 demonstrate its discriminative power and capacity to maintain solid precision and recall across classes.
The differences between the macro- and weighted-average metrics (e.g., Macro F1-Score of 0.74 versus Weighted F1-Score of 0.90) highlight that certain less frequent classes may not be identified as accurately as more common ones. The Log Loss of 0.35 and Brier Score of 0.15 indicate relatively good calibration, suggesting that the model’s probability estimates are reasonably aligned with observed outcomes. Additionally, a Hamming Loss of 0.08 and a Specificity (Macro Avg) of 0.89 confirm that the model makes relatively few incorrect predictions and can identify negative instances correctly. These findings suggest that, while SecureFalcon performs admirably, further improvements—such as targeted data augmentation or calibration adjustments—could enhance its robustness across all vulnerability categories.
VI-D Comparison of SecureFalcon with LLM, ML, and DL models